Which is an equation of a line that has a slope of -1/2 and contains the point (2,3)?
Answer by checkley71(8403)

Y=mX+b WHERE m=SLOPE & b= Y INTERCEPT.
3=-1/2*2+b
3=-1+b
b=3+1
b=4 THE Y INTERCEPT.
THUS THE EQUATION IS:
Y=-X/2+4